# FisherRF: Active View Selection and Uncertainty Quantification for Radiance Fields using Fisher Information

### v1

@@ 3.5 Pixel-wise Uncertainty with Volumetric Rendering
Previously, we discussed quantifying the uncertainty for any camera views. Our model can also be extended to obtain pixel-wise uncertainties. As we have discussed in Sec. 3.3, we can approximate the uncertainty on each parameter with the diagonal elements of the hessian matrix ${\mathbf{H}}^{\prime\prime}[{\bf y}^{acq}_{i}|{\bf x}^{acq}_{i},{\bf w}^{*}]$. In recent neural rendering models [13, 29], the parameters directly correspond to a spatial location in the scene. Therefore, we can compute the uncertainty in our rendered pixels by examining the diagonal Hessians along the casted ray for volumetric rendering
@@
where ${\bf G}_{n}$ is the submatrix of ${\mathbf{H}}^{\prime\prime}[{\bf w}^{*}|D^{train}]$ containing the rows and columns that correspond to parameters at location $n$. Please note this is a relative uncertainty on the rendered pixels that treat each location equally. If we wish to estimate absolute uncertainties on predictions like depth maps, we need to denormalize the uncertainty in each term by its depth $d_{n}$.
@@ 4 Experiments
In this Section, we present the empirical evaluation of our approach. First, we introduce the details of our implementation in (Subsection 4.1). Then, we focus on the experiments of active view selection and compare our method quantitatively and qualitatively against the previous method. Finally, we present more results of our method on uncertainty quantification. results (Subsection 4.3).
4.1 Implementation Details
Our method can be applied to various kinds of neural rendering models. We implement the computation of Fisher Information on Plenoxels [29] and 3D Gaussian Splatting [13] with customized CUDA kernels.
As shown in Eq. 10, the diagonal Hessian matrixes can be implemented as efficiently as a back-propagation. Therefore, our customized CUDA kernel that computes diagonal Hessians for 3D Gaussian Splatting achieved more than 70 fps on a Nvidia RTX3090 GPU. The log-prior regularizer $\lambda$ in Eq. 11 is $10^{-6}$ across all experiments. The supplementary materials and code release will present other hyper-parameters and implementation details.
@@ 4.2 Active View Selection
We conducted extensive experiments on view selections to demonstrate that our expected information gain could help the model find the next best views. Here, we first introduce the dataset we use and detailed experimental settings. Then, we compare our method with random baselines and previous state-of-the-art [25] quantitatively and qualitatively.
@@
Our approach is extensively evaluated on two common benchmark datasets: Blender Dataset [22] and the challenging real-world Mip-NeRF360 dataset [3]. The Blender dataset comprises eight synthetic objects with intricate geometry and realistic non-Lambertian materials. Each scene in this dataset includes 100 training and 200 test views, all with a resolution of $800\times 800$. Our method uses the 100 training views as a candidate pool to select training views, and we evaluate all the models on the full 200 views test set. We use the default training configuration as in 3D Gaussian and Plenoxels for this dataset. Mip-NeRF360 [3] is a real-world $360^{\circ}$ dataset captured for nine different scenes. It has been widely used as a quality benchmark for novel view synthesis models [4, 23]. We train our models at the resolution of $1066\times 1600$ following 3D Gaussian Splatting [13].
@@
We quantitatively and qualitatively compare our method against the current state-the-art ActiveNeRF [25] and random selection baseline. To make a fair comparison with ActiveNeRF [25], we re-implemented a similar variance estimation algorithm in 3D Gaussian Splatting and Plenoxel in CUDA. We assign each 3D Gaussian (or grid cell in Plenoxels) a variance parameter and use volumetric rendering to render a variance map. More details about our re-implementation can be found in the supplementary materials.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
PSNR/SSIM 400 Steps 8000 Steps 350k Steps
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: A comparative study on the effects of the number of training views and view selection strategies. The top renderings are from the random baseline, and the bottom are from our method. Both are implemented with 3D Gaussian Splatting. We further extend our experiments on the Blender dataset from selecting 20 views to selecting 80 views. We continue training over the model by selecting one new view every 100 epochs. Although the effects of view selection vanish after the model is trained with a sufficient number of views, our method is still better than the random baseline as it was less impacted by the degeneration and local minima when the model was only trained with a limited number of views at the early stage.
ActiveNeRF Random Ours
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 6: Qualitative Results on Blender Dataset with 20 Training Views All the methods are implemented on 3D Gaussian Splatting and compared in the same condition except for different training views selected by different methods. As can be seen, ActiveNeRF suffered from significant artifacts at a large scale, and the random baseline also experienced blurry details when zooming in. Our method, however, could select the training views that are most necessary for the current model, thus preventing catastrophic degeneration.
ActiveNeRF Random Ours
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 7: Qualitative Results on Blender Dataset with 10 Training Views We also qualitatively compare the method with ActiveNeRF and random baseline implemented on top of 3D Gaussian Splatting. By selecting the training views with maximum information gain, our model could still help the 3D Gaussian Splatting to get impressive views under this challenging setting.
@@
ActiveNeRF* 26.240 0.8560 0.1240
Plenoxel + Random 23.242 0.8617 0.1582
Plenoxel + ActiveNeRF 23.522 0.8573 0.1499
Plenoxel + ActiveNeRF†) 23.147 0.8571 0.1478
3D Gaussian + Random 28.732 0.9389 0.0534
3D Gaussian + Random† 27.135 0.9267 0.0651
3D Gaussian + ActiveNeRF 25.854 0.9157 0.0766
3D Gaussian + ActiveNeRF† 27.326 0.9116 0.0756
Plenoxel + Ours 24.513 0.8759 0.1568
Plenoxel + Ours† 24.212 0.8782 0.1389
3D Gaussian + Ours 29.525 0.9436 0.0431
3D Gaussian + Ours† 29.094 0.9379 0.0531
@@
Plenoxel + Random 19.950 0.8124 0.2329
Plenoxel + ActiveNeRF 19.770 0.8044 0.2098
3D Gaussian + Random 20.670 0.8242 0.2049
3D Gaussian + ActiveNeRF 22.979 0.8756 0.1109
Plenoxel + Ours 19.770 0.8044 0.2098
3D Gaussian + Ours 23.681 0.8831 0.1021
Table 2: Active View Selections on Blender Dataset with only 10 views The best, second, and third results are highlighted in red, orange, and yellow, respectively. Our view selection algorithm could select necessary views when the number of training views is extremely limited.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
3D Gaussian + Random† 19.542 0.5684 0.3759
3D Gaussian + ActiveNeRF 17.889 0.5326 0.4142
3D Gaussian + ActiveNeRF† 18.303 0.5391 0.4059
3D Gaussian + Ours 20.351 0.6010 0.3608
3D Gaussian + Ours† 20.568 0.6078 0.3647
[…]
Statue $\downarrow$ Africa $\downarrow$ Torch $\downarrow$ Basket $\downarrow$ Average $\downarrow$
@@
Ours 0.26 0.30 0.33 0.19 0.27
Table 4: Uncertainty Estimation on LF Dataset from $360^{\circ}$ views Numbers are AUSE, the lower the better. The best results are highlighted in red, and the second-best results are in orange color. Our model outperformed previous state-of-the-art on most scenes by a large margin.
The quantitative results of active view selections can be found in Table. 1 and Table. 3. As can be seen, our method achieved better results across different metrics and datasets. We also compare our method qualitatively on the Blender and Mip-NeRF360 datasets in Fig. 4 and Fig. 6. Our sequential selection variant achieved better results on the Synthetic dataset because our batch acquisition algorithm is a greedy approximation of sequential view selection. However, the benefit of sequential view selection vanishes in the challenging real-world dataset because selecting necessary views at the early stage of training is crucial to prevent degeneration or local minima. This is further supported by our visualizations in Fig. 4 and Fig. 5, where we show the baseline models exhibited obvious artifacts due to insufficient regularizations across different training views. This suggests that batch view selection is preferable in real-world applications as it enables view planning without significant compromise on performance. Furthermore, we experiment with our model with the challenging ten views selection task on Blender Dataset. Each model is initialized with the same random seed and two uniform initial views. Each new view is added after every 100 epochs till the model has ten training views. The quantitative and qualitative results are in Table. 2 and Fig. 7. Again, our method selects necessary views given the extremely limited observations and preserves fine details of reconstructed objects.
@@ 4.3 Uncertainty Quantification
As discussed in Sec. 3.5, our model can be extended to compute pixel-wise uncertainties on training views. Following previous methods on uncertainty estimation [32, 31], we evaluate our method on the Light Field (LF) Dataset [39] using the Area Under Sparsification Error (AUSE) metric. The pixels are filtered twice, once using the absolute error with ground truth depth and once using the uncertainty. The difference in the mean absolute error on the remaining pixels between the two sparsification processes produces two different error curvatures, where the area between those two curvatures is the AUSE, which evaluates the correlation between uncertainties and the predicted error. A low AUSE indicates our model is confident in the correctly estimated depths and could predict a high uncertainty in the regions where we are likely to have larger errors.
As 3D Gaussian Splatting is not designed for forward-facing scenes, we re-run the CF-NeRF on $360^{\circ}$ views by using every 10th view for each scene in the LF dataset as a fixed training set and every 16th view as the test set. As seen in Table 4, our model exhibited better results than the previous state-of-the-art CF-NeRF [32]. More quantitative results and visualizations on uncertainty estimation can be found in our supplementary materials.

### v2

@@ 7       $S_{b}\leftarrow S_{b-1}\cup\{i\}$ ;
3.4 Active Mapping with 3D Gaussian Splatting
Figure 2: An Illustration of Our Active Mapping System Given RGBD captures, we first reconstruct the environment using 3D Gaussians as representation. Afterward, we select a set of goal points from map frontiers and plan the shortest path to each goal. The EIG is computed for each path, and we choose the one with the highest EIG to continue exploration.
With the batch selection algorithm discussed in Sec. 3.3, our method can be extended as an active mapping algorithm. By maximizing the EIG from candidate trajectories $\{\mathcal{P}_{j}\}$, we can determine the most informative path for exploration:
$\operatorname*{arg\,max}_{\mathcal{P}_{j}}\operatorname{tr}\left(\left(\sum_{{\bf x}^{acq}_{i}\in\mathcal{P}_{j}}{\mathbf{H}}^{\prime\prime}[{\bf y}^{acq}_{i}|{\bf x}^{acq}_{i},{\bf w}^{*}]\right)\;{\mathbf{H}}^{\prime\prime}[{\bf w}^{*}|D^{train}]^{-1}\right).$ (9)
We build our active mapping system on top of existing passive mapping architecture SplaTAM [24] that uses 3D Gaussians as basic data representation. We employ frontier-based exploration [64] to propose path candidates and sample cameras along the path concerning the mobility of the agent. The pipeline is illustrated in Fig. 2. Our path selection algorithm is compatible with other active mapping systems and can be used as an extension to select the most informative path.
@@ 3.5 Pixel-wise Uncertainty with Volumetric Rendering
Previously, we discussed quantifying the Information Gain for any camera views. Our model can also be adapted to obtain pixel-wise uncertainties. As we have discussed in Sec. 3.2, we can approximate the uncertainty on each parameter with the diagonal elements of the hessian matrix ${\mathbf{H}}^{\prime\prime}[{\bf y}^{acq}_{i}|{\bf x}^{acq}_{i},{\bf w}^{*}]$. In recent Radiance Field models [25, 52], the parameters directly correspond to a spatial location in the scene. Therefore, we can compute the uncertainty in our rendered pixels by examining the diagonal Hessians along the casted ray for volumetric rendering
@@
where ${\bf G}_{n}$ is the submatrix of ${\mathbf{H}}^{\prime\prime}[{\bf w}^{*}|D^{train}]$ containing the rows and columns that correspond to parameters at location $n$. Please note that this is a relative uncertainty derived from the observed information on the model parameters, which does not involve metric scales. If we wish to estimate absolute uncertainties on predictions like depth maps, we need to denormalize the uncertainty in each term by its depth $d_{n}$.
@@ 4 Experiments
In this Section, we present the empirical evaluation of our approach. The running time analysis for our customized CUDA kernel and other implementation details can be found in the supplementary. We first focus on the experiments of active view selection and compare our method quantitatively and qualitatively against the previous method (Subsection 4.1). We further present our results on active mapping (Subsection 4.2). Finally, we present more results of our method on uncertainty quantification results (Subsection 4.3).
@@ 4.1 Active View Selection
We conducted extensive experiments on view selections to demonstrate that our expected information gain could help the model find the next best views. Here, we first introduce the dataset we use and detailed experimental settings. Then, we compare our method with random baselines and previous state-of-the-art ActiveNeRF [43] quantitatively and qualitatively.
@@
Our approach is extensively evaluated on two common benchmark datasets: Blender Dataset [38] and the challenging real-world Mip-NeRF360 dataset [5]. The Blender dataset comprises eight synthetic objects with intricate geometry and realistic non-Lambertian materials. Each scene in this dataset includes 100 training and 200 test views, all with a resolution of $800\times 800$. Our method uses the 100 training views as a candidate pool to select training views, and we evaluate all the models on the full 200-view test set. We use the default training configuration as in 3D Gaussian and Plenoxels for this dataset. Mip-NeRF360 [5] is a real-world $360^{\circ}$ dataset captured for nine different scenes. It has been widely used as a quality benchmark for novel view synthesis models [6, 39]. We train our models at the resolution of $1066\times 1600$ following 3D Gaussian Splatting [25].
@@
We quantitatively and qualitatively compare our method against the current state-the-art ActiveNeRF [43] and random selection baseline. To make a fair comparison with ActiveNeRF [43], We re-implemented a similar variance estimation algorithm in 3D Gaussian Splatting and Plenoxels in CUDA. We assign each 3D Gaussian (or grid cell in Plenoxels) a variance parameter and use volumetric rendering to render a variance map. The supplementary materials provide more details about our re-implementation. To show the difference between EIG and uncertainty, we adapt the uncertainty estimation method BayesRays [16] into the view selection algorithm by directly using the rendered uncertainty value on each candidate view as a metric for view selection.
ACNeRF Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Random Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ours Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GT Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
20 Training Views 10 Training Views
ActiveNeRF Random BayesRays Ours ActiveNeRF Random BayesRays Ours
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Qualitative Results on Blender Dataset with 20 and 10 Training Views. All the methods are implemented on 3D Gaussian Splatting and compared in the same training configuration except for different training views selected by different methods.
@@ 20 Training Views 10 Training Views
Table 1: Active View Selections on Blender Dataset The best, second, and third results are highlighted in red, orange, and yellow, respectively.  *: Numbers are taken directly from ActiveNeRF [43] as reference.
@@
ActiveNeRF* 26.240 0.856 0.124
Plenoxels + Random 23.906 0.863 0.174
Plenoxels + ActiveNeRF 23.522 0.857 0.150
Nerfacto + Random 22.614 0.890 0.111
Nerfacto + BayesRays 22.504 0.887 0.115
3D Gaussian + Random 28.732 0.939 0.053
3D Gaussian + ActiveNeRF 26.610 0.905 0.081
Plenoxel + Ours 24.513 0.876 0.157
3D Gaussian + Ours 29.525 0.944 0.043
Table 2: Active Batch View Selections on Blender Dataset Each time, 4 views are selected based on different view selection methods. *: Numbers are taken directly from ActiveNeRF [43] as reference.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
ActiveNeRF* 26.240 0.856 0.124
Plenoxels + Random 23.242 0.862 0.158
Plenoxels + ActiveNeRF 23.147 0.857 0.148
Nerfacto + Random 21.622 0.875 0.137
Nerfacto + BayesRays 22.160 0.878 0.127
3D Gaussian + Random 27.135 0.927 0.065
3D Gaussian + ActiveNeRF 27.326 0.912 0.076
Plenoxels + Ours 24.212 0.878 0.139
3D Gaussian + Ours 29.094 0.938 0.053
Table 3: Active View Selections on Blender Dataset with only 10 views Our view selection algorithm could select necessary views when the number of training views is extremely limited.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Nerfacto + Random 16.430 0.809 0.254
Nerfacto + BayesRays 14.950 0.798 0.282
Plenoxels + Random 19.950 0.812 0.233
Plenoxels + ActiveNeRF 19.770 0.804 0.210
3D Gaussian + Random 22.493 0.873 0.112
3D Gaussian + ActiveNeRF 22.979 0.876 0.111
Plenoxels + Ours 20.670 0.824 0.205
3D Gaussian + Ours 23.681 0.883 0.102
@@
3D Gaussian + Random 17.914 0.564 0.430
3D Gaussian + Random† 19.542 0.568 0.376
3D Gaussian + ActiveNeRF 17.889 0.533 0.414
3D Gaussian + ActiveNeRF† 18.303 0.539 0.406
3D Gaussian + Ours 20.351 0.601 0.361
3D Gaussian + Ours† 20.568 0.608 0.365
Table 5: Scene Coverage on Gibson and MP3D dataset Our method outperformed previous methods by a large margin when evaluating the coverage metric. Numbers for other methods are taken from Active Nerual Mapping [67].
Method Gibson MP3D
Comp. (%) $\uparrow$ Comp. (cm) $\downarrow$ Comp. (%$\uparrow$ Comp. (cm) $\downarrow$
Random 45.80 34.48 45.67 26.53
FBE 68.91 14.42 71.18 9.78
UPEN 63.30 21.09 69.06 10.60
OccAnt 61.88 23.25 71.72 9.40
Active Neural Mapping 80.45 7.44 73.15 9.11
Ours 92.89 5.64 89.41 2.91
The quantitative results of active view selections can be found in Table. 2, Table. 2 and Table. 5. As can be seen, our method achieved better results across different metrics and datasets. We also compare our method qualitatively on the Blender and Mip-NeRF360 datasets in Fig. 3 and Fig. 4. Our method demonstrates superior results in both single-view selection and batch-view selection. Moreover, our method achieved better performance improvements compared to the random backbone compared to BayesRays. This further renders the importance of using EIG rather than uncertainty as the metric for view selection. Furthermore, we experimented with our model with the challenging ten-view selection task on the Blender Dataset. Each model is initialized with the same random seed and two uniform initial views. Each new view is added after every 100 epochs till the model has ten training views. The quantitative and qualitative results are in Table. 3 and Fig. 4. Again, our method selects necessary views given the extremely limited observations and preserves fine details of reconstructed objects.
4.2 Active Mapping
We experimented with our active mapping system on room-scale scene dataset Matterport3D [7] and Gibson [63] through the Habitat Simulator [53, 47]. Our active mapping system takes posed RGB-D images as input and uses the same configurations for the simulator as previous approaches [67, 15]. We turn our 3D Gaussian representation into a point cloud using only the mean $\mathbf{\mu}\in R^{3}$ of each 3D Gaussian as the previous evaluation benchmark is designed for point cloud or 3D mesh. We use the following evaluation metric to evaluate the quality of our active mapping system.
Comp. The completeness metric quantifies the completeness of the active exploration algorithm in 3D space by computing the per-point nearest distance between points sampled from the ground truth mesh and the final reconstruction of the scene. We calculate both the percentage of points within a 5cm threshold (Comp. (%)) and the average nearest distance (Comp. ($cm$)).
PSNR & Depth MAE We uniformly sample 1000 navigable poses in the scene and render color and depth images at the sampled poses and compute average PSNR and Mean Absolute Error (MAE) for rendered color and depth images, respectively.
Ground Truth Active-INR Ours
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 5: Qualitative Comparisons on Active Mapping We compare our method against the ground truth mesh and Active Neural Mapping. Our method exhibited better details and coverage. Our method can successfully produce detailed and high-fidelity reconstruction for the rooms compared to the previous method.
Table 6: Evaluation for Render Quality on Gibson and MP3D We also compare our method with previous method on the rendering quality for both color and depth reconstruction.
Method Gibson MP3D
PSNR $\uparrow$ Dpeth MAE (m) $\downarrow$ PSNR $\uparrow$ Depth MAE (m) $\downarrow$
FBE 19.35 0.1751 16.68 0.3627
UPEN 21.13 0.1893 18.40 0.2690
Active Neural Mapping 19.64 0.2943 17.16 0.4621
Ours 22.58 0.0924 19.96 0.1667
Table 7: Uncertainty Estimation on LF Dataset Numbers are AUSE, the lower the better. The best results are highlighted in red, and the second-best results are in orange color.
Method Statue $\downarrow$ Africa $\downarrow$ Torch $\downarrow$ Basket $\downarrow$ Average $\downarrow$
@@ 4.2 Active Mapping
ActiveNeRF 0.40 0.39 0.21 0.34 0.33
BayesRays 0.17 0.27 0.22 0.28 0.23
Ours 0.21 0.26 0.24 0.18 0.22
We compare with previous state-of-the-art in Table. 5 and Figure. 5. Our active mapping system outperformed previous methods by a large margin on both geometry and rendering quality. Notably, we are the first active mapping system to capture fine-grained details and high-quality textures, thanks to our view selection objectives, which aim to improve rendering quality when measuring the EIG. In Figure 5, we demonstrate the qualitative comparison between our method and Active-INR [67]. In Table 5, our method outperforms Active-INR by a large margin in Comp. metric, showing the superiority of the active mapping algorithm. To compare the rendering quality with previous methods, we train a 3D Gaussian Splatting model using the trajectory from UPEN [15] and Active-INR [67] and evaluate the render quality at the same holdout set. The result is summarized in Table 7. Our method efficiently selects the most informative path and outperforms other methods when using the same backbone model. We also notice that UPEN is worse than Active Neural Mapping in Table 5 but better than the other in 7 when evaluated using the 3D Gaussian Splatting model. We ascribe this gap in UPEN to the difference in the scene representations and the evaluation protocol.
@@ 4.3 Uncertainty Quantification
CF-NeRF BayesRays Ours
Uncertainty Depth Err. Uncertainty Depth Err. Uncertainty Depth Err.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: Visualizations for Uncertainty on Depth Prediction the LF Dataset We compare the per-pixel uncertainty maps from CF-NeRF [55], BayesRays [16], and ours. Our per-pixel uncertainty maps demonstrate a strong correlation with the actual depth error.
As discussed in Sec. 3.5, our model can be extended to compute pixel-wise uncertainties on training views. Following previous methods on uncertainty estimation [55, 56], we evaluate our method on the Light Field (LF) Dataset [69] using the Area Under Sparsification Error (AUSE) metric. The pixels are filtered twice, once using the absolute error with ground truth depth and once using the uncertainty. The difference in the mean absolute error on the remaining pixels between the two sparsification processes produces two different error curvatures, where the area between those two curvatures is the AUSE, which evaluates the correlation between uncertainties and the predicted error. A low AUSE indicates our model is confident in the correctly estimated depths and could predict a high uncertainty in the regions where we are likely to have larger errors. As seen in Table 7 and Fig. 6, our model exhibited a better correlation between our uncertainty estimation and depth error than the previous state-of-the-art CF-NeRF [55]. More quantitative results and visualizations on uncertainty estimation can be found in our supplementary materials.